OpenStudy (anonymous):

Solve. –13 = –7 + v A. –20 B. –6 C. 6 D. 20

OpenStudy (igreen):

-13 = -7 + v Add 7 to both sides: -13 + 7 = -7 + 7 + v -7 + 7 cancels out leaving 'v' by itself. -13 + 7 = v Add: v = -6

OpenStudy (anonymous):

so it would be a negative?

OpenStudy (igreen):

Yes, it would.

OpenStudy (anonymous):

ok im going to close this ?